# Projected and Solvable Topological Heavy Fermion Model of Twisted Bilayer Graphene

### v1

@@ II The topological heavy fermion model
$H^{(c,\eta)}(\mathbf{k})=\begin{pmatrix}0_{2\times}&v_{\star}(\eta k_{x}\sigma_{0}+ik_{y}\sigma_{z})\\ v_{\star}(\eta k_{x}\sigma_{0}-ik_{y}\sigma_{z})&M\sigma_{x}\end{pmatrix},\qquad H^{(cf,\eta)}(\mathbf{k})=\begin{pmatrix}\gamma\sigma_{0}+v_{\star}^{\prime}(\eta k_{x}\sigma_{x}+k_{y}\sigma_{y})\\ 0_{2\times 2}\end{pmatrix}\ .$ (4)
@@
where $U_{1}$ is the interaction strength and $:f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}:=f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}-\frac{1}{2}$. Other terms include the density-density interactions between $f$- and $c$-electrons $\hat{H}_{W}$, density-density interactions between nearest neighbors of $f$-electrons $H_{U_{2}}$, a Hund’s exchange between $f$- and $c$-electrons $\hat{H}_{J}$, and Coulomb repulsions between $c$ electrons $\hat{H}_{V}$. All the interaction parameters are also determinate functions of the $\theta$ and $w_{0}/w_{1}$ Călugăru et al. (2023). The single-particle and interaction parameters at the magic angle are given in supplementary tables S4 and S6 of Ref. Song and Bernevig (2022). The full interacting theory, with terms beyond $U_{1}$ treated at the mean-field level, is detailed in an extensive DMFT-Iterative Perturbation Theory–Hubbard-I manuscript Călugăru and others .
@@ III Wavefunction of the topological flat band
where $\mathcal{N}_{\mathbf{k}}=1+\sum_{\mathbf{G}}\frac{(\gamma/v_{\star})^{2}}{|\mathbf{k}+\mathbf{G}|^{2}}e^{-\lambda^{2}|\mathbf{k}+\mathbf{G}|^{2}}$ is a normalization factor. To derive the analytical formula of topological flat-band wavefunction, we omit all $|\mathbf{G}|>0$ components. Using Eqs. 2 and 3, we obtain the following Bloch function with Chern number $+1$ in valley $+$
@@
Using Eq. 11, we could obtain the following expression for the concentrated Berry curvature of the Chern $+1$ band in valley $+$
@@
where $\chi_{l\beta}$ is a spinor in valley and sub-lattice space, $\mathbf{r}\!=\!(x,y)$ is restricted in the first moiré unit cell, and $\mathbf{k}$ is restricted in the first moiré Brillouin zone. Here $\delta$ controls the spread of the concentrated charge in AA stacking regions, and $s$ controls the spread of Berry curvature in momentum space. When $\mathbf{r}$ ($\mathbf{k}$) exceeds the first unit cell (Brillouin zone), periodic conditions $\psi_{l\beta;\mathbf{k}+\mathbf{G}}^{\rm NW}(\mathbf{r})=\psi_{l\beta;\mathbf{k}}^{\rm NW}(\mathbf{r})$, $\psi_{l\beta;\mathbf{k}+\mathbf{G}}^{\rm NW}(\mathbf{r}+\mathbf{R})=e^{\mathrm{i\>\!}\mathbf{k}\cdot\mathbf{R}}\psi^{\rm NW}_{l\beta;\mathbf{k}}(\mathbf{r})$ are enforced. A subtlety of Eq. 13 is that $|\psi^{\rm NW}_{l\beta;\mathbf{k}}(\mathbf{r})|^{2}$ exhibits artificial discontinuities in both momentum and real spaces. In Ref. Ledwith et al. (2024) $\delta$ and $s$ are treated as fitting parameters to minimize the error in the intra-Chern-band form factors. The inter-Chern-band form factors are omitted. This automatically leads to an artificial chiral symmetry as well as a U(4)$\times$×U(4) symmetry in the projected interaction, as explained in Ref. Bernevig et al. (2021d), even though the parameters $\delta,s$ are fitted using realistic parameters without chiral symmetry ($w_{0}\neq 0$). We termed such projected interaction as in a quasi-chiral limit which has also been discussed in Ref. Song and Bernevig (2022).
@@ IV.1 Self-energy and Hubbard-I approximation
[…]
@@
[…]
@@
This describes an atomic problem where $f$ electrons of different sites are decoupled. We use $\nu_{f}$ to characterize the filling of the $f$ electrons. The atomic action $S_{at}$ can be solved exactly which yields the following self-energy in the low-energy limit for the $f$ electrons at the integer filling ($1/\beta\ll U_{1}$ălugăru and others
@@
The single-particle excitation of the system can be obtained by finding the poles of Green’s function ($[G^{\alpha\eta s}_{flat}(i\omega\rightarrow E_{\mathbf{k}},\mathbf{k})]^{-1}=0$) which gives (at $\nu=0$)
@@ IV.1.1 Single-particle spectrum in the Hubbard-I approximation across all interaction regimes
In this section, we discuss the single-particle spectrum in the Hubbard-I approximation in detail. We first utilize the idea introduced in Section VII.1 and map the interacting Green’s function to an effective non-interacting Green’s function. We note that the self-energy in Eq. 24 can be written as
[…]
@@
[…]
@@
±$\pm\bigg{(}\frac{M}{2}\pm\sqrt{\tilde{V}^{2}|\mathbf{k}|^{2}+M^{2}/4}\bigg{)}$ (43)
@@ IV.1.2 Comparison of the Hubbard-I spectrum in two limits
±$\pm\sqrt{\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}-\sqrt{\bigg{[}\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}\bigg{]}^{2}-\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}}{4}}}$ (46)
@@
±$\pm\sqrt{\frac{U_{1}^{2}}{4\gamma^{2}+U_{1}^{2}}}|v_{\star}\mathbf{k}|$ (47)
@@ IV.2 Effective coupling between flavor moments
[…]
@@
The diagonal components of $S$ with a value of $1/2$ correspond to the flavors that are filled, while those with a value of $-1/2$ correspond to the flavors that are empty. This gives $\text{Tr}[O(\mathbf{R})]=\text{Tr}[S]=\nu$. The decomposition in Eq. 70 gives
@@
Only terms with $n_{1}\in 2\mathbb{Z}+1,n_{2}\in 2\mathbb{Z}+1$ generate a coupling between moments ($O(\mathbf{R})$) between different sites. Therefore, we focus on the terms with $11+k\in 2\mathbb{Z}+1,n-1-k\in 2\mathbb{Z}+$ 1 in  Eq. 73. This also implies that $n\in 2\mathbb{Z},k\in 2\mathbb{Z}$. We let $n=2m,k=2s$, and the corresponding contributions are
@@ VI Quantitative aspects of the THF model
In this limit, the Berry curvature becomes highly concentrated and is effectively described by a $\delta$-function (Eq. 12). However, such a curvature can only be realized in a system with band touching. This implies that the gap between the remote bands and the dispersive bands vanishes, making it necessarily smaller than any finite $U_{1}$, which itself must be larger than the bandwidth of the narrow bands. This can be shown on general grounds (see Section VIII) and can also be observed within the THF model. If we fix the bandwidth of the $c$-electrons ($|v_{\star}\Lambda_{c}|$), then taking the limit $\gamma/|v_{\star}\Lambda_{c}|\to 0$ necessarily leads to $\gamma\to 0$, indicating that the gap between the remote and flat bands disappears in the heavy-fermion model. Thus, in this limit, due to band touching, one cannot simply focus on (or project onto) the flat bands while neglecting the remote bands. However, the heavy-fermion model, which explicitly incorporates both flat and remote bands, remains a valid framework for studying both regimes.
@@ Acknowledgements.
VII.1 Appendix: Mapping an interacting Green’s function to a non-interacting Green’s function
@@
[…]
where $\Sigma_{ij}(i\omega)$ is the self-energy. We can utilize the spectral decomposition of the self-energy which reads
[…]

### v2

@@ II The topological heavy fermion model
$H^{(c,\eta)}(\mathbf{k})=\begin{pmatrix}0_{2\times}&v_{\star}(\eta k_{x}\sigma_{0}+ik_{y}\sigma_{z})\\ v_{\star}(\eta k_{x}\sigma_{0}-ik_{y}\sigma_{z})&M\sigma_{x}\end{pmatrix},\qquad H^{(cf,\eta)}(\mathbf{k})=\begin{pmatrix}\gamma\sigma_{0}+v_{\star}^{\prime}(\eta k_{x}\sigma_{x}+k_{y}\sigma_{y})\\ 0_{2\times 2}\end{pmatrix}\$. (4)
@@
where $U_{1}$ is the interaction strength and $f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}:=f_{\mathbf{R},\alpha,\eta,s}^{\dagger}f_{\mathbf{R},\alpha,\eta,s}-\frac{1}{2}$. Other terms include the density-density interactions between $f$- and $c$-electrons $\hat{H}_{W}$, density-density interactions between nearest neighbors of $f$-electrons $H_{U_{2}}$, a Hund’s exchange between $f$- and $c$-electrons $\hat{H}_{J}$, and Coulomb repulsions between $c$ electrons $\hat{H}_{V}$. All the interaction parameters are also determinate functions of the $\theta$ and $w_{0}/w_{1}$ Călugăru et al. (2023). The single-particle and interaction parameters at the magic angle are given in supplementary tables S4 and S6 of Ref. Song and Bernevig (2022). The full interacting theory, with terms beyond $U_{1}$ treated at the mean-field level, is detailed in an extensive DMFT-Iterative Perturbation Theory–Hubbard-I manuscript Călugăru and others .
@@ III Wavefunction of the topological flat band
where $\mathcal{N}_{\mathbf{k}}=1+\sum_{\mathbf{G}}\frac{(\gamma/v_{\star})^{2}}{|\mathbf{k}+\mathbf{G}|^{2}}e^{-\lambda^{2}|\mathbf{k}+\mathbf{G}|^{2}}$ is a normalization factor. To derive the analytical formula of topological flat-band wavefunction, we omit all $|\mathbf{G}|>0$ components. Using Eqs. 2 and 3, we obtain the following Bloch function with Chern number +1+1 in valley ++
@@
Using Eq. 11, we could obtain the following expression for the concentrated Berry curvature of the Chern +1+1 band in valley ++
@@
[…]
@@ IV.1 Self-energy and Hubbard-I approximation
$+\frac{U_{1}}{2}\int_{0}^{\beta}d\tau\sum_{\mathbf{R},\alpha\eta s,\alpha^{\prime}\eta^{\prime}s^{\prime}}\left(f_{\mathbf{R},\alpha\eta s}^{\dagger}(\tau)f_{\mathbf{R},\alpha\eta s}(\tau)-\frac{1}{2}\right)\left(f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}^{\dagger}(\tau)f_{\mathbf{R},\alpha^{\prime}\eta^{\prime}s^{\prime}}(\tau)-\frac{1}{2}\right)$ (20)
@@
$=$ $\frac{\gamma^{2}}{v_{\star}^{2}\Lambda^{2}}\frac{2\pi}{\Omega_{M}}\int_{\epsilon<|v_{\star}\Lambda_{c}|}d\epsilon\left[\frac{\epsilon/2}{\mathrm{i\>\!}\omega+\epsilon_{c}-\epsilon}+\frac{\epsilon/2}{\mathrm{i\>\!}\omega+\epsilon_{c}+\epsilon}\right]J_{0}\left(\frac{\epsilon\cdot|\mathbf{R}-\mathbf{R}^{\prime}|}{|v_{\star}|}\right)$ (21)
@@
This describes an atomic problem where $f$ electrons of different sites are decoupled. We use $\nu_{f}$ to characterize the filling of the $f$ electrons. The atomic action $S_{at}$ can be solved exactly which yields the following self-energy in the low-energy limit for the $f$ electrons at the integer filling ($1/\beta\ll U_{1}$Călugăru and others
@@
The single-particle excitation of the system can be obtained by finding the poles of Green’s function ($G^{\alpha\eta s}_{flat}(i\omega\rightarrow E_{\mathbf{k}},\mathbf{k})]^{-1}=0$) which gives (at $\nu=0$)
@@ IV.1.1 Single-particle spectrum in the Hubbard-I approximation across all interaction regimes
In this section, we discuss the single-particle spectrum in the Hubbard-I approximation in detail. We first utilize the idea introduced in Section VIII.3 and map the interacting Green’s function to an effective non-interacting Green’s function. We note that the self-energy in Eq. 24 can be written as
${\Sigma}_{f}\left(i\omega\right)=\int_{-\infty}^{\infty}\frac{\rho^{{\Sigma}_{f}}(\epsilon)}{i\omega-\epsilon}d\epsilon,\quad\rho^{{\Sigma}_{f}}(\epsilon)=\frac{U_{1}^{2}(4-\nu_{f})(\nu_{f}+4)}{64}\delta\left(\epsilon-U_{1}\nu_{f}/8\right)$ (34)
@@
$+\sum_{\mathbf{k},\alpha\eta s}U_{1}\nu_{f}/8a_{\mathbf{k},\alpha\eta s}^{\dagger}a_{\mathbf{k},\alpha\eta s}+\sum_{\mathbf{k},\alpha\eta s}\frac{U_{1}\sqrt{(4-\nu_{f})(\nu_{f}+4)}}{8}\left(a_{\mathbf{k},\alpha\eta s}^{\dagger}{f}_{\mathbf{k},\alpha\eta s}+\text{h.c.}\right).$ (35)
@@
$\pm\bigg(\frac{M}{2}\pm\sqrt{\tilde{V}^{2}|\mathbf{k}|^{2}+M^{2}/4}\bigg)$ (43)
@@ IV.1.2 Comparison of the Hubbard-I spectrum in two limits
$\pm\sqrt{\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}-\sqrt{\bigg[\frac{\gamma^{2}+|v_{\star}\mathbf{k}|^{2}+(\frac{U_{1}}{2})^{2}}{2}\bigg]^{2}-\frac{U_{1}^{2}|v_{\star}\mathbf{k}|^{2}}{4}}}$ (46)
@@
$\pm\sqrt{\frac{U_{1}^{2}}{4\gamma^{2}+U_{1}^{2}}}|v_{\star}\mathbf{k}|$ (47)
@@ IV.2 Effective coupling between flavor moments
$g_{0}(\tau-\tau^{\prime})=-\langle T_{\tau}d_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau)d^{\dagger}_{\mathbf{k},(-1)^{\alpha+1},\eta s}(\tau^{\prime})\rangle,\quad g_{0}(i\omega)=\int_{0}^{\beta}g_{0}(\tau)d\tau$ (59)
@@
The diagonal components of $S$ with a value of $1/2$ correspond to the flavors that are filled, while those with a value of 1/2-1/2 correspond to the flavors that are empty. This gives $\text{Tr}[O(\mathbf{R})]=\text{Tr}[S]=\nu$. The decomposition in Eq. 70 gives
@@
Only terms with $n_{1}\in 2\mathbb{Z}+1,n_{2}\in 2\mathbb{Z}+1$ generate a coupling between moments ($O(\mathbf{R})$) between different sites. Therefore, we focus on the terms with $1+k\in 2\mathbb{Z}+1,n-1-k\in 2\mathbb{Z}+1$ in  Eq. 73. This also implies that $n\in 2\mathbb{Z},k\in 2\mathbb{Z}$. We let $n=2m,k=2s$, and the corresponding contributions are
@@ VI Quantitative aspects of the THF model
In this limit, the Berry curvature becomes highly concentrated and is effectively described by a $\delta$-function (Eq. 12). However, such a curvature can only be realized in a system with band touching. This implies that the gap between the remote bands and the dispersive bands vanishes, making it necessarily smaller than any finite $U_{1}$, which itself must be larger than the bandwidth of the narrow bands. This can be shown on general grounds (see Section IX) and can also be observed within the THF model. If we fix the bandwidth of the $c$-electrons ($|v_{\star}\Lambda_{c}|$), then taking the limit $\gamma/|v_{\star}\Lambda_{c}|\to 0$ necessarily leads to $\gamma\to 0$, indicating that the gap between the remote and flat bands disappears in the heavy-fermion model. Thus, in this limit, due to band touching, one cannot simply focus on (or project onto) the flat bands while neglecting the remote bands. However, the heavy-fermion model, which explicitly incorporates both flat and remote bands, remains a valid framework for studying both regimes.
@@ Acknowledgements.
VIII Appendix: Mapping an interacting Green’s function to a non-interacting Green’s function
@@
$G_{ij}(i\omega)=\int_{0}^{\beta}G_{ij}(\tau)e^{i\omega_{n}\tau}d\tau=\bigg[i\omega\mathbb{I}-t-\Sigma(i\omega)\bigg]_{ij}^{-1}$ (87)
where $\Sigma_{ij}(i\omega)$ is the self-energy.
VIII.1 Properties of Green’s function
We first discuss the properties of the Green’s function. For a generic electron operator $\gamma_{i}$ with flavor index $i$ (momentum, orbital, spin, sublattice, etc.). We define the Green’s function as
$G_{ij}(\tau)=-\langle T_{\tau}\gamma_{i}(\tau)\gamma^{\dagger}_{j}(0)\rangle$ (88)
We assume the eigenstates and eigenenergy of the Hamiltonians are $|n\rangle,E_{n}$ respectively with
$H|n\rangle=E_{n}|n\rangle$ (89)
Then from the definition of Green’s function we have (for $\beta>\tau\geq 0$)
$G_{ij}(\tau)=-\frac{1}{Z}\sum_{n,m}e^{-(\beta-\tau)E_{n}-\tau E_{m}}\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle,\quad Z=\sum_{n}e^{-\beta E_{n}}$ (90)
Performing Fourier transformation, we could obtain the Green’s function in the Matsubara frequency
$G_{ij}(i\omega)=\int_{0}^{\beta}G_{ij}(\tau)e^{i\omega_{n}\tau}d\tau=\frac{1}{Z}\sum_{n,m}\frac{1}{i\omega-E_{m}+E_{n}}\bigg(e^{-\beta E_{m}}+e^{-\beta E_{n}}\bigg)\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle$ (91)
By performing analytical continuation, the Green’s function on the entire complex plane reads
$G_{ij}(z)=\int_{0}^{\beta}G_{ij}(\tau)e^{i\omega_{n}\tau}d\tau=\frac{1}{Z}\sum_{n,m}\frac{1}{z-E_{m}+E_{n}}\bigg(e^{-\beta E_{m}}+e^{-\beta E_{n}}\bigg)\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle$ (92)
The asymptotic behaviors of Green’s function at large $|z|$ is
$G_{ij}(z)\sim$ $\frac{1}{Z}\sum_{n,m}\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle\bigg(e^{-\beta E_{m}}+e^{-\beta E_{n}}\bigg)\frac{1}{z}\bigg[1+\frac{E_{m}-E_{n}}{z}+O(1/z^{2})\bigg]$
$=$ $\sum_{n}\frac{\langle n|\gamma_{i}\gamma_{j}^{\dagger}+\gamma_{j}^{\dagger}\gamma_{i}|n\rangle e^{-\beta E_{n}}}{Z}\frac{1}{z}+\frac{N_{ij}}{z^{2}}+O(1/z^{3})$
$=\frac{\delta_{i,j}}{z}+\frac{N_{ij}}{z^{2}}+O(1/z^{3})$ (93)
where
$N_{ij}=\frac{1}{Z}\sum_{n,m}\sum_{n,m}\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle(e^{-\beta E_{m}}+e^{-\beta E_{n}})(E_{m}-E_{n})$ (94)
We can also observe that $N_{ij}=N_{ji}^{*}$ indicating $N$ is an Hermitian matrix.
The spectral functions are
$\rho_{ij}(\omega)=\frac{1}{\pi}\text{Im}[G_{ij}(\omega-i0^{+})]=\frac{1}{Z}\sum_{n,m}\delta(\omega-E_{m}+E_{n})\bigg(e^{-\beta E_{m}}+e^{-\beta E_{n}}\bigg)\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle$ (95)
From the definition of the spectral function, we can observe that it is Hermitian with
$[\rho_{ij}(\omega)]^{*}=\rho_{ji}(\omega)$ (96)
In addition, we also find
$G_{ij}(z)=\int_{-\infty}^{\infty}\frac{\rho_{ij}(\omega)}{z-\omega}d\omega$ (97)
which corresponds to the spectral representation of Green’s function. The spectral function also follows the sum rule
$\int_{-\infty}^{\infty}\rho_{ij}(\omega)=\frac{1}{Z}\sum_{n,m}\bigg(e^{-\beta E_{m}}+e^{-\beta E_{n}}\bigg)\langle n|\gamma_{i}|m\rangle\langle m|\gamma_{j}^{\dagger}|n\rangle$
$=$ $\frac{1}{Z}\bigg[\sum_{m}e^{-\beta E_{m}}\langle m|\gamma_{j}^{\dagger}\gamma_{i}|m\rangle+\sum_{n}e^{-\beta E_{n}}\langle n|\gamma_{i}\gamma_{j}^{\dagger}|n\rangle\bigg]=\frac{1}{Z}\sum_{m}\bigg[e^{-\beta E_{m}}\langle m|\gamma_{j}^{\dagger}\gamma_{i}+\gamma_{i}^{\dagger}\gamma_{j}|m\rangle\bigg]$
$=$ $\delta_{i,j}\frac{1}{Z}\sum_{m}e^{-\beta E_{m}}$
$=$ $\delta_{i,j}$ (98)
VIII.2 Properties of self-energy
In this section, we discuss the properties of the self-energy. From the definition, it behaves as
$\Sigma_{ij}(z)=z\delta_{i,j}-[H_{0}]_{ij}-G_{ij}^{-1}(z)$ (99)
In the large $|z|\gg 1$ limit we have
$\Sigma_{ij}(z)\sim z\delta_{i,j}-[H_{0}]_{ij}-z\delta_{i,j}+N_{ij}+O(1/z^{2})\sim-[H_{0}]_{ij}+N_{ij}+O(1/z)$ (100)
Therefore $\Sigma_{ij}(z)$ may not decay to zero at $|z|\rightarrow\infty$. However, one could separate the constant contributions from the self-energy and let
$\Sigma_{ij}(z)=\delta\Sigma_{ij}(z)+N_{ij}-[H_{0}]_{ij}$ (101)
where $\delta\Sigma_{ij}(z)$ is an analytical function and decay as $1/|z|$. Therefore, $\delta\Sigma(z)$ follows Kramer-Kronig relations.
VIII.3 Mapping an interacting Green’s function to a non-interacting Green’s function
The Green’s function is then
$G_{ij}(i\omega)=\int_{0}^{\beta}G_{ij}(\tau)e^{i\omega_{n}\tau}d\tau=\bigg[i\omega\mathbb{I}-t-\Sigma(i\omega)\bigg]_{ij}^{-1}$ (103)
where $\Sigma_{ij}(i\omega)$ is the self-energy. We could separate the self-energy into two parts
$\Sigma(i\omega)=\delta\Sigma(i\omega)+N$ (104)
with $N=\lim_{|\omega|\rightarrow\infty}\Sigma(\omega)$. We can utilize the spectral decomposition of the self-energy which reads
$\delta\Sigma_{ij}(i\omega_{n})=\int_{-\infty}^{\infty}\frac{\rho^{\Sigma}_{ij}(\epsilon)}{i\omega_{n}-\epsilon}d\epsilon$ (105)